What is the mass of the precipitate formed when 50 mL of 20% solution of AgNO(w/v) is mixed with 50 mL of 6% (w/v) NaCl solution?

Grade:11

1 Answers

Vikas TU
3 years ago
Dear student 
No of moles of AgNo3 present is 
(50*16.9) /(100*169.8) = 0.0497
Mole of NaCl present is 
(50*5.8)/(100*58.5) = 0.0496
Thus, both AgNO3 and NaCl are present in the stoichiometric quantity.
The number of moles of AgCl formed is 0.0496. 
The mass of AgCl formed is 0.0496×143.3=7 g.